# CVPT: Cross-Attention help Visual Prompt Tuning adapt visual task

### v1

@@ 4.3 Ablation Studies
The impact of the location of the Cross-Attention (CA). We conducted experiments with the following five positions to explore the optimal deployment of CA, and the results of the experiments are displayed in Table.5:
@@
Figure 5: (a) The deployments of cross-attention in ViT. Five possible positions can be inserted. Our final deployments are in dark blue. (b) Performance comparisons of different deployments of cross-attention.
We can see that inserting in prompt tokens after self-attention (SA) is the best way to perform. However, if a slight performance decrease is acceptable, we can choose position 2 to insert in parallel to improve the efficiency of the operation (this improvement is also slight).
The impact of weight-sharing between CA and SA. We set CA to be learnable (without weight-sharing) and frozen (with weight-sharing) respectively to investigate the impact of weight-sharing. The results on VTAB-1K and FGVC are shown in Table.5 below:
@@
Setting Learnable Para(M) VTAB-1K FGVC
Nat. Spe. Str. Avg.
learnable CA 28.4 80.1 84.8 57.8 74.2 89.4
frozen CA 0.08 80.1 84.4 57.8 74.1 90.3
We find that setting CA to tunable adds a significant number of parameters, substantially increasing computational overhead. Despite the slight performance gain it brings on VTAB-1K, it lags behind the frozen CA substantially in FGVC. Therefore, We believe that the parameters of SA are valuable for guiding the fine-tuning of CA. Especially, when dealing with a complete dataset of a certain size, such as FGVC, the weight-sharing mechanism can better utilize the pre-trained capabilities of the model, thereby improving performance.
@@ 5 Conclusion
In this paper, we explore the current mainstream prompt-based method VPT deeply and analyze the reasons why it performs poorly. Consequently, we propose a simple and effective PEFT method, CVPT, which introduces the cross-attention module to compute the cross-attention between the prompt tokens and embedded tokens thus instructing the model’s fine-tuning. What more, the weights of cross-attention are come from self-attention, avoiding introducing an enormous number of additional trainable parameters and achieving better performance. We conducted extensive experiments on 25 datasets, and the results demonstrate that CVPT achieves SOTA performance. Additionally, we conducted extensive ablation experiments on CVPT, demonstrating the impact of introducing cross-attention and weight-sharing, as well as its efficiency and performance improvements over VPT. We hope our work will inspire prompt-based PEFT methods in the future. One limitation of our work is that CVPT does not explore new strategies for the initialization of prompt tokens. In VPT, the author made a complete comparison of different initialization methods. In our work, we take the same strategy with VPT. However, we still think the optimized specific initialization method is better than the general methods VPT used. Besides, this initialization will also help us understand how prompts help the model’s fine-tuning.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: See abstract, introduction, method and experiments.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: See conclusion. We think a good strategy which we don’t mention in this paper can help improving the performance based on our work.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: We don’t think our work involves that.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: See experimental settings.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: See Abstract.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: See experimental settings.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: We follow the previous works.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: See experimental settings.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We don’t think our works in relation to this.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: We don’t think our work involves that.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: We don’t think our work involves that.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We used publicly available datasets whose licenses allow research usage.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: We don’t think our work involves that.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: We don’t think our work involves that.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: We don’t think our work involves that.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.

### v2

@@ 5.4 Ablation Studies
The impact of the location of the Cross-Attention (CA). We conducted experiments with the following five positions to explore the optimal deployment of CA, and the results of the experiments are displayed in Fig.5:
It can be observed that positions near the self-attention (SA) module (1-3) outperform those near the MLP block (4-5), and we attribute this to the more effective adaptation of the rich, contextual features generated by SA. Prior work [36] also demonstrated that inserting adapters near SA yields better results. Among all options, inserting prompt tokens after SA (position 3) achieves the best performance. Notably, position 2 offers a compelling trade-off, enabling parallel insertion for improved efficiency with only a negligible drop in performance.
@@
Figure 5: (a) The deployments of cross-attention in ViT. Our final deployments are in dark blue. (b) Performance comparisons of different deployments.
The impact of weight-sharing between CA and SA. Weight-sharing can be viewed as an initialization strategy for CA. We introduce random init for comparison and include linear probing (fine-tuning only the classifier head) to demonstrate the effectiveness of introduced prompts and frozen CA. Results with 10 prompts are shown in Table.6:
@@
Method Nat. Spe. Str. Avg FGVC Param(M)
Linear probing 68.9 77.2 26.8 57.6 79.3 0
Random init+Leanable CA 80.0 84.9 57.2 74.0 89.5 28.4
Weight Sharing+Leanable CA 79.8 85.3 58.8 74.6 89.5 28.4
Random init+Frozen CA 72.8 81.8 36.6 63.7 86.0 0.09
Weight Sharing+Frozen CA 80.0 84.5 57.5 74.0 89.3 0.09
After introducing weight-sharing, a frozen module significantly outperforms one with random initialization. Notably, its performance is comparable to that of a fully learnable CA, despite requiring substantially fewer parameters. This strongly suggests that the parameters inherited from SA provide a powerful inductive bias, effectively guiding the CA module.
While a larger parameter count can increase model capacity, it does not guarantee better performance and may lead to overfitting. The frozen CA, with its smaller and well-initialized parameter space, is therefore easier to optimize and converges much faster than the fully learnable version, proving to be a more efficient approach.
@@ 6 Conclusion
In the current field of visual fine-tuning, many researchers overlook prompts in favor of adapters due to their strong performance. The few prompt-based derived works do not realize the drawbacks of combining prompts with embedded tokens, continuing to use the method from VPT. In light of this, we thoroughly analyzed the shortcomings of such deployment and proposed CVPT. Its advantages are as follows: 1) It decouples prompts from self-attention and uses cross-attention to establish a connection with embedded tokens. 2) It employs weight-sharing to avoid the large number of learnable parameters introduced by cross-attention. Additionally, we conducted extensive experiments on CVPT, demonstrating its efficiency and performance improvements over VPT and the effectiveness of cross-attention and weight-sharing. Therefore, we prove that prompt-based methods can perform comparably to advanced adapter methods in the visual fine-tuning domain. We hope our work will inspire prompt-based PEFT methods in the future.
The limitation of our work is that we do not propose new strategies for the init of prompt tokens. We have made some attempts and noticed that other studies have been introduced [27, 50]. However, neither our attempts nor those reported elsewhere have shown substantial effects. Therefore, we follow the same with VPT. We think the exploration of initialization will help us understand how prompts help the model’s fine-tuning.
7 Acknowledgement
This work was supported in part by the National Natural Science Foundation of China under Grant 62133005, Grant 62293512 and Grant 62293515, the Special funding support for the construction of innovative provinces in Hunan Province under Grant 2021GK1010, the Central Guidance Fund Project for Hunan Science and Technology Development under Grant 2023ZYT003-1.